Find the Maclaurin series for the given function. - 3x ΟΑ. Σε Ο8. Σ - οο. Στο M8 M8 M8 M8
Answer #1

Solutions - Given f (c) = e 3 We know that Maclausin series for fx is given be f(x) = f(0) + f(a) x + f/o) x 2 43%)23 - sinceinc. f (o)= E) - Putting the values foron 2.@undD in D, we have fle = ctz 1 + (D33+ ***7480P - - -+-):, , = 1 + 1) 3) ac,